Mr. Smith receives a monthly income of $3,500 from his pension plan. It takes two days for the cheque to be mailed to him, one day for him to deposit the cheque at his bank, and five days for the bank to process the cheque and put the money in his bank account. What is the average daily float? Use 30 days per month.
Answer #1

Total float = 2+1+5 = 8 days

Average daily float = Check amount*Float days/ Total days in month

=3500*8/30

=$933.33